# Active Learning of Model Discrepancy with Bayesian Experimental Design

### v1

@@ 2.3 Ensemble-based Indicator for Data Informativeness
where $G$ denotes a forward map from unknown parameters $\bm{\theta}$ to the data $\mathbf{y}$, and $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\bm{\Gamma})$ denotes the data measurement noises and is often assumed to be with a zero-mean Gaussian distribution. In this work, the forward map $G$ corresponds to solving the modeled system in Eq. (2) and observing the solution according to a given design $\mathbf{d}$, and $\bm{\theta}$ corresponds to the unknown parameters $\bm{\theta}_{\text{NN}}$ of the model discrepancy term.
@@
$\begin{gathered}D_{\textrm{KL}}(p(\bm{\theta}|\mathbf{y})||p(\bm{\theta}))\approx\tilde{D}_{\textrm{KL}}(\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}||\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1})=\\ \frac{1}{2}\left[\text{tr}\left(\left(\Sigma^{\bm{\theta}\bm{\theta}}_{K}\right)^{-1}\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)-d_{\bm{\theta}}+\ln\left(\frac{\det\Sigma^{\bm{\theta}\bm{\theta}}_{K}}{\det\Sigma^{\bm{\theta}\bm{\theta}}_{0}}\right)\right.\left.+\left(\bar{\bm{\theta}}_{K}-\bar{\bm{\theta}}_{0}\right)^{\top}\left(\Sigma^{\bm{\theta}\bm{\theta}}_{K}\right)^{-1}\left(\bar{\bm{\theta}}_{K}-\bar{\bm{\theta}}_{0}\right)\right],\end{gathered}$ (18)
@@
The ensemble-based approximation of KLD in Eq. (18) can be used to evaluate the informativeness of the data $\mathbf{y}$ from a design $\mathbf{d}$. In this work, we focus on applying the approximated KLD for the high-dimensional parameters $\bm{\theta}_{\text{NN}}$ in Eq. (2) and determine whether to perform the calibration of the model discrepancy based on the optimal design and the corresponding measurements from the standard BED of the relatively low-dimensional parameters $\bm{\theta}_{\mathcal{G}}$. In practice, a few iterations of EKI updating are sufficient to evaluate the data informativeness, which can effectively avoid performing calibration of the model discrepancies on less informative or even misleading data that could further lead to larger model discrepancies and more biases in the estimation of $\bm{\theta}_{\mathcal{G}}$ in Eq. (2). More detailed algorithm to exploit the ensemble-based approximation of KLD to quantify the data informativeness is presented in A for the iterative approach introduced in Section 2.2.
@@ 3 Numerical Results
$\frac{\partial\mathbf{u}(\mathbf{z},t;\bm{\theta})}{\partial t}=\nabla^{2}\mathbf{u}-\mathbf{v}(t)\cdot\nabla\mathbf{u}+S(\mathbf{z},t;\bm{\theta}),~{}~{}~{}\mathbf{z}\in[z_{L},z_{R}]^{2},~{}~{}t>0,$ (20)
where $\mathbf{v}=\{v_{x},v_{y}\}\in\mathbb{R}^{2}$ is a time-dependent convection velocity, $S$ denotes the source term with some parameters $\bm{\theta}$. In this work, the true system has an exponentially decay source term in the following form with the parameters $\bm{\theta}=\{\theta_{x},\theta_{y},\theta_{h},\theta_{s}\}\in\mathbb{R}^{4}$:
@@
where $\theta_{x}$ and $\theta_{y}$ denote the source location, $\theta_{h}$ and $\theta_{s}$ represent the source width and source strength. The initial condition is $\mathbf{u}(\mathbf{z},0;\bm{\theta})=\mathbf{0}$, and a homogeneous Neumann boundary condition is imposed for all sides of the square domain.
@@
Figure 2: Concentration value at different times in a convection-diffusion field. The numerical simulation is performed in a larger domain ($[-2,3]^{2}$) and presented in a smaller one ($[0,1]^{2}$) to emphasize the areas close to the source location. Stages 1-5 correspond to $0.05,0.10,0.15,0.20,0.25$ time units.
The design $\mathbf{d}=\{d_{x},d_{y},d_{t}\}$ in this problem refers to the spatiotemporal coordinate to measure the concentration value. More specifically, each design involves the measurement of just one point in the domain. The spatial coordinates of the initial design start at $\{d_{x}^{0},d_{y}^{0}\}=(0.5,0.5)$ and will gradually move around to other locations. The differences of spatial coordinates between two consecutive designs (i.e., $d_{x}^{i+1}-d_{x}^{i}$ and $d_{y}^{i+1}-d_{y}^{i}$) are constrained by the interval $[-0.2,0.2]^{2}$. The temporal coordinates of the measurement are set with a fixed time step $\Delta t=0.05$ time unit. At each time step, the posterior distribution from the previous stage serves as the prior for a new stage of BED as illustrated in Fig. 1. The current optimal design also serves as the starting point for the next stage of BED.
@@ 3.1 Validation of Iterative Approach for Parametric Model Error
In this section, the performance of our proposed iterative approach is validated by an example with parametric model error. The setup of this example is designed in such a way that the full BED remains feasible. Compared with the results of full BED approach, we demonstrate that the proposed iterative approach can efficiently and robustly calibrate the parametric model error. More specifically, we set up an example with parametric model error where the true form of the source function in Eq. (21) is known but the value of $\theta_{s}$ is set incorrectly, e.g., due to the lack of knowledge on the strength of the source term. In addition, the source location $(\theta_{x},\theta_{y})$ are unknown and to be determined by BED in the presence of parametric model error $\theta_{s}$. The true values are $\theta_{x}=0.45$, $\theta_{y}=0.25$, and $\theta_{s}=2$. The advection velocity is known and set as $v_{x}=v_{y}=20t$.
@@ 3.1.1 Standard Approach on Parametric Model Error
We first consider joint probability distribution of $\{\theta_{x},\theta_{y},\theta_{s}\}$ and perform a five-stage standard sequential BED as the benchmark results. The five stages correspond to 0.05, 0.10, 0.15, 0.20, and 0.25 time units. For the posterior distribution at each stage, we first present the marginal distribution of $\theta_{s}$, followed by the two-dimensional conditional distribution of $\theta_{x}$ and $\theta_{y}$, conditioned on the value of $\theta_{s}$ corresponding to the highest marginal probability density. The results in Fig. 3 demonstrate that the standard approach successfully identifies the true values of the parameters. For $\theta_{s}$, a peak of probability density gradually forms around the true value of 2 in its marginal distribution, indicating increased confidence in the estimation. In the two-dimensional distribution of $\theta_{x}$ and $\theta_{y}$, the high probability region also aligns well with the true source location.
@@ 3.1.2 Hybrid Approach on Parametric Model Error
At the first stage of sequential BED, we employ a uniform prior distribution of $\{\theta_{x},\theta_{y}\}$ over the range $[0,1]^{2}$, discretized into a $51\times 51$ grid of points, and set an initial value of $\theta_{s}=3$. In each stage, BED is performed to infer the posterior distribution of $\{\theta_{x},\theta_{y}\}$, and the optimal design is determined accordingly based on the information gain introduced in Section 2.1. The optimal design subsequently provides additional information that facilitates the calibration of the model error $\theta_{s}$. In terms of the gradient-based optimization of model error, we have tested both automatic differentiation and adjoint method to acquire the gradient of $\mathrm{d}L/\mathrm{d}\theta_{s}$ and achieved comparable results. For simplicity and clarity, we only show the results from the automatic differentiation approach, where we have developed a numerical solver based on JAX-CFD [64, 65] to facilitate the extraction of gradient information.
@@
Figure 4 shows the evolution of the posterior distribution for $\{\theta_{x},\theta_{y}\}$ across five stages. The high probability regions gradually converge to the true source location, suggesting an effective correction of parametric model error $\theta_{s}$. The comparison between Figs 3 and 4 confirms that the proposed hybrid approach achieves similar results to standard BED for the joint distribution of $\{\theta_{x},\theta_{y},\theta_{s}\}$ but with lower computational cost.
@@
Figure 4: Posterior results of OED parameter $\{\theta_{x},\theta_{y}\}$ via the hybrid approach on parametric error. The $\{\theta_{x},\theta_{y}\}$ space is $[0,1]^{2}$ as before, with a zoomed-in view $[0.2,0.6]\times[0.1,0.5]$ to highlight detailed behaviors.
@@
We further investigate the results of corrected parametric model error $\theta_{s}$. As shown in Fig. 5(a), the proposed hybrid approach identifies the true value of $\theta^{\dagger}_{s}=2$. It is worth noting that the optimization process of $\theta_{s}$ tends to converge to an incorrect value between iterations 450 and 600 (stage 4), mainly due to the poor data quality associated with the design at that stage for model error correction. To address this issue, we employ the approximated KLD defined in Eq. (18) to evaluate the data informativeness. The poor-quality design at stage 4 in Fig. 5(a) not only renders a bias in the corrected value for the parametric model error $\theta_{s}$ but also leads to a slower rate of variance reduction during EKI iterations in Fig. 5(b), which implies a smaller approximated KLD between the prior and posterior ensemble of $\theta_{s}$. The comparison in Fig. 5(b) confirms that the performance of the ensemble-based approximated KLD as an indicator to evaluate data informativeness for the calibration of parametric model error.
@@ 3.2 Hybrid Approach on Structural Errors
which is different from the exponentially decaying source term as defined in Eq. (21) for the true system. Without addressing this model structural error, the inference of source location through standard BED yields biased results. In this work, a neural network $\text{NN}(z_{x},z_{y},\theta_{x},\theta_{y};\bm{\theta_{\text{NN}}})$ is employed to characterize the model structural error. More specifically, we utilize a fully connected neural network consisting of 37 parameters, a configuration that introduces a higher-dimensional parameter space than the example of parametric model error in Section 3.1.
@@
We employ a uniform prior distribution over the range $[0,1]^{2}$ for $\{\theta_{x},\theta_{y}\}$, discretized into a $51\times 51$ grid, with the true value of the source location set at $\{0.25,0.25\}$. The values of $\{\theta_{s}^{\dagger},\theta_{h}^{\dagger}\}$ are set as $\{2,0.05\}$ in both the true system and the modeled one. The advection velocity is known and set as $v_{x}=v_{y}=50t$.
@@
We first present the results of standard sequential BED for physics-based parameters updated without correcting the structural error. It is expected that model discrepancy introduces bias in both the model forecast and likelihood values, resulting in an incorrect posterior distribution, i.e., the high probability area does not cover the true source location in Fig. 6(a). Fig. 6(b) shows that the posterior distribution can be improved by applying the proposed hybrid framework to correct the model error, and the high probability areas gradually converge to the true source location from stage 1 to stage 3. The deviation of high probability areas from the true source location in stages 4 and 5 is mainly caused by the uninformative design and data used for model error correction at stage 3. These design and data, optimized for physical parameters at stage 3, introduce biases into the calibrated model, which propagate into the Bayesian updating of physical parameters in subsequent stages.
@@
Figure 7 shows that a more informative design and the corresponding measurement data (adopted at stage 3) for the model error correction significantly improve the posterior distributions of inferred source location at stages 4 and 5. We present the results using the ensemble-based approximated KLD to assess the data informativeness of a given design in Fig. 8. Two designs are selected for comparison: (i) the design identified by BED for posterior estimation at stage 3, represented by the orange dot in Fig. 7, and (ii) a modified design represented by the blue triangle in Fig. 8. With the untrained model at stage 3 as a baseline, we first generate an ensemble of its parameters and apply EKI to update the ensemble for 10 iterations using the two designs and corresponding measurement data. The ensemble-based approximated KLD between the updated and initial ensembles at each EKI iteration in Fig. 8(b) shows a faster increase with the good design, which confirms that the ensemble-based indicator proposed in Section 2.3 is effective of assessing data informativeness. The model error updated by the good design also shows less deviation from the true discrepancy in Fig. 8(a). Considering that the measurement data at a few design points are not sufficient to correct the model error term across the whole domain, some remaining differences in Fig. 8(a) between the true source term and the modeled one are expected. Even with those differences in the modeled source term, especially at a smaller distance from the source location, the posterior distributions in Fig. 7 still demonstrate a good estimation of the inferred source location.
@@
MSE $=\frac{1}{N}\sum_{z_{x},z_{y}}(\mathbf{u}(z_{x},z_{y})-\mathbf{u}^{\dagger}(z_{x},z_{y}))^{2},$ (23)

### v2

@@ 2.3 Ensemble-based Indicator for Data Informativeness
where $G$ denotes a forward map from unknown parameters $\bm{\theta}$ to the data $\mathbf{y}$, and $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\bm{\Gamma})$ denotes the data measurement noises and is often assumed to be with a zero-mean Gaussian distribution. In this work, the forward map $G$ corresponds to a composition of solving the modeled system in Eq. (2) and obtaining the observational data according to a given design $\mathbf{d}$. The parameters $\bm{\theta}$ only account for the unknown parameters $\bm{\theta}_{\text{NN}}$ of the model discrepancy term, while the parameters $\bm{\theta}_{\mathcal{G}}$ are considered to be fixed.
@@
$\begin{gathered}D_{\textrm{KL}}(p(\bm{\theta}|\mathbf{y})||p(\bm{\theta}))\approx\tilde{D}_{\textrm{KL}}(\{\bm{\theta}_{K}^{(j)}\}^{J}_{j=1}||\{\bm{\theta}_{0}^{(j)}\}^{J}_{j=1})=\\ \frac{1}{2}\left[\text{tr}\left(\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\Sigma^{\bm{\theta}\bm{\theta}}_{K}\right)-d_{\bm{\theta}}+\ln\left(\frac{\det\Sigma^{\bm{\theta}\bm{\theta}}_{0}}{\det\Sigma^{\bm{\theta}\bm{\theta}}_{K}}\right)\right.\left.+\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)^{\top}\left(\Sigma^{\bm{\theta}\bm{\theta}}_{0}\right)^{-1}\left(\bar{\bm{\theta}}_{0}-\bar{\bm{\theta}}_{K}\right)\right],\end{gathered}$ (18)
@@
The ensemble-based approximation of KLD in Eq. (18) can be used to evaluate the informativeness of the data $\mathbf{y}$ from a design $\mathbf{d}$. In this work, we focus on applying the approximated KLD for the high-dimensional parameters $\bm{\theta}_{\text{NN}}$ in Eq. (2) and determine whether to perform the calibration of the model discrepancy based on the optimal design and the corresponding measurements from the standard BED of the relatively low-dimensional parameters $\bm{\theta}_{\mathcal{G}}$. In practice, a few iterations of EKI updating are sufficient to evaluate the data informativeness, which can effectively avoid performing calibration of the model discrepancies on less informative or even misleading data that could further lead to larger model discrepancies and more biases in the estimation of $\bm{\theta}_{\mathcal{G}}$ in Eq. (2). The selection of these informative design points is iterative: starting from an initial or random candidate, its KLD value and the approximate gradient are determined. New points are then chosen along the KLD ascent direction, and this is repeated until a predefined budget is met. More detailed algorithm to exploit the ensemble-based approximation of KLD to quantify the data informativeness is presented in Algorithm 1 for the iterative approach introduced in Section 2.2.
Algorithm 1 Active Learning of Model Discrepancy with BED
for i=1,2,,Ni=1,2,...,N do $\triangleright$ Iterating stage
  $\mathbf{d}\leftarrow\operatorname*{arg\,max}_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[U(\bm{\theta}_{\mathcal{G}},\mathbf{y},\mathbf{d};\bm{\theta}_{\textrm{NN}})]$ $\triangleright$ $\bm{\theta}_{\textrm{NN}}$ is fixed
  $\mathbf{y}\leftarrow\mathbf{u}(\mathbf{d})+\bm{\epsilon}$ $\triangleright$ Measurement at optimal design
  $\bm{\theta}_{\mathcal{G}}\leftarrow\bm{\theta}_{\mathcal{G}}|\mathbf{y}$ $\triangleright$ Bayesian update
  $\bm{\theta}_{\mathcal{G}}^{*}\leftarrow\operatorname*{arg\,max}_{\bm{\theta}_{\mathcal{G}}}p(\bm{\theta}_{\mathcal{G}};\bm{\theta}_{\text{NN}})$ $\triangleright$ MAP
   $\bm{\theta}_{\text{NN}}\leftarrow\operatorname*{arg\,max}_{\bm{\theta}_{\text{NN}}}L(\bm{\theta}_{\text{NN}};\bm{\theta}_{\mathcal{G}}^{*},\mathbf{y},\mathbf{d})$ $\triangleright$ $\bm{\theta}_{\mathcal{G}}^{*}$ is fixed
  end if
end for
@@ 3 Numerical Results
$\frac{\partial\mathbf{u}(\mathbf{z},t;\bm{\theta})}{\partial t}=D\nabla^{2}\mathbf{u}-\mathbf{v}(t)\cdot\nabla\mathbf{u}+S(\mathbf{z},t;\bm{\theta}),~~~\mathbf{z}\in[z_{L},z_{R}]^{2},~~t>0,$ (20)
where $D$ is the diffusion coefficient (assumed to be known with a true value of $1$), $\mathbf{v}=\{v_{x},v_{y}\}\subseteq\mathbb{R}^{2}$ is a time-dependent convection velocity, $S$ denotes the source term with some parameters $\bm{\theta}$. In this work, the true system has an exponentially decay source term in the following form with the parameters $\bm{\theta}=\{\theta_{x},\theta_{y},\theta_{h},\theta_{s}\}\in\mathbb{R}^{4}$:
@@
where $\theta_{x}$ and $\theta_{y}$ denote the source location, $\theta_{h}$ and $\theta_{s}$ represent the source width and source strength. The initial condition is $\mathbf{u}(\mathbf{z},0;\bm{\theta})=\mathbf{0}$, and a homogeneous Neumann boundary condition is imposed for all sides of the square domain. The convection term is discretized using the second-order total variation diminishing Van Leer scheme, while the diffusion term is treated implicitly via a fast diagonalization method. Time integration follows a semi-implicit scheme: convection is updated explicitly using the forward Euler method, and diffusion is solved implicitly to ensure numerical stability. The time step is chosen as 0.0005 to satisfy the Courant-Friedrichs-Lewy condition imposed by the explicit convection term. The Peclet number in our setup is approximately 0.2, indicating a diffusion-dominated regime where no additional stabilization is required.
Note that in this work, the inversion is performed over a finite-dimensional parameter set $\bm{\theta}$ which governs either the source term (e.g., source location, width, and strength) or the parametric form of the convection velocity (e.g., a time-dependent velocity with unknown coefficient $k_{y}$ in $v_{y}=k_{y}t$). The state variable $\mathbf{u}$ not directly inferred but is obtained by solving the governing PDE given $\bm{\theta}$ and $k_{y}$.
@@
Figure 2: Concentration value at different times in a convection-diffusion field. The numerical simulation is performed in a larger domain ($2,3]^{2}$) and presented in a smaller one ($0,1]^{2}$) to emphasize the areas close to the source location. Stages 1-5 correspond to $0.05,0.10,0.15,0.20,0.25$ time units.
The design $\mathbf{d}=\{d_{x},d_{y},d_{t}\}$ in this problem refers to the spatiotemporal coordinate to measure the concentration value. More specifically, each design involves the measurement of just one point in the domain. The spatial coordinates of the initial design start at $\{d_{x}^{0},d_{y}^{0}\}=(0.5,0.5)$ and will gradually move around to other locations. The differences of spatial coordinates between two consecutive designs (i.e., $d_{x}^{i+1}-d_{x}^{i}$ and $d_{y}^{i+1}-d_{y}^{i}$) are constrained by the interval $0.2,0.2]^{2}$. Unlike traditional sensor selection problems where sensors can be placed freely across the entire domain, our setting assumes either a single movable sensor or a localized sensor array that relocates sequentially. The local constraint was initially introduced to ensure that each design decision meaningfully influences the feasible region for future designs, creating a dependency between stages. This local movement constraint is also motivated by a realistic setting where the sensor is movable but has limited relocation capability due to physical or temporal restrictions. The temporal coordinates of the measurement are set with a fixed time step $\Delta t=0.05$ time units. At each time step, the posterior distribution from the previous stage serves as the prior for a new stage of BED as illustrated in Fig. 1. The current optimal design also serves as the starting point for the next stage of BED.
@@
  • 4.
    We also examine another more challenging scenario in which the convection coefficient $k_{y}$ in the velocity field $v_{y}=k_{y}t$ is also unknown, an extension beyond the source-term functional-form discrepancy in Section 3.2. We compare two strategies: (i) treating $v_{y}$ as a fixed and incorrect constant, and (ii) jointly calibrating $k_{y}$ together with the neural network parameters $\bm{\theta}_{\text{NN}}$. This example underscores the performance of our framework for an ill-posed problem of unknown parameters estimation. Detailed results can be found in Section 3.3.
In addition to the numerical example studied in this section, we also test the performance of our method on another numerical example of an acoustic inverse problem and present some key results in C.
@@ 3.1 Validation of Iterative Approach for Parametric Model Error
In this section, the performance of our proposed iterative approach is validated by an example with parametric model error. The setup of this example is designed in such a way that the full BED remains feasible. Compared with the results of full BED approach, we demonstrate that the proposed iterative approach can efficiently and robustly calibrate the parametric model error. More specifically, we set up an example with parametric model error where the true form of the source function in Eq. (21) is known but the value of $\theta_{s}$ is set incorrectly, e.g., due to the lack of knowledge on the strength of the source term. In addition, the source location $\{\theta_{x},\theta_{y}\}$ is unknown and to be determined by BED in the presence of parametric model error $\theta_{s}$. The true values are $\theta_{x}=0.45$, $\theta_{y}=0.25$, and $\theta_{s}=2$. The advection velocity is known and set as $v_{x}=v_{y}=20t$.
@@ 3.1.1 Standard Approach on Parametric Model Error
We first consider joint probability distribution of $\{\theta_{x},\theta_{y},\theta_{s}\}$ and perform a five-stage standard sequential BED as the benchmark results. The prior distribution of $\{\theta_{x},\theta_{y},\theta_{s}\}$ is assumed to be uniform on the domain $0,1]\times[0,1]\times[1,4]$, discretized into a $51\times 51\times 31$ grid. The five stages correspond to 0.05, 0.10, 0.15, 0.20, and 0.25 time units. For the posterior distribution at each stage, we first present the marginal distribution of $\theta_{s}$, followed by the two-dimensional conditional distribution of $\theta_{x}$ and $\theta_{y}$, conditioned on the value of $\theta_{s}$ corresponding to the highest marginal probability density. The results in Fig. 3 demonstrate that the standard approach successfully identifies the true values of the parameters. For $\theta_{s}$, a peak of probability density gradually forms around the true value of 2 in its marginal distribution, indicating increased confidence in the estimation. In the two-dimensional distribution of $\theta_{x}$ and $\theta_{y}$, the high probability region also aligns well with the true source location.
@@ 3.1.2 Hybrid Approach on Parametric Model Error
At the first stage of sequential BED, we employ a uniform prior distribution of $\{\theta_{x},\theta_{y}\}$ over the range $0,1]^{2}$, discretized into a $51\times 51$ grid of points, and set an initial value of $\theta_{s}=3$. In each stage, BED is performed to infer the posterior distribution of $\{\theta_{x},\theta_{y}\}$, and the optimal design is determined accordingly based on the information gain introduced in Section 2.1. The optimal design subsequently provides additional information that facilitates the calibration of the model error $\theta_{s}$. In terms of the gradient-based optimization of model error, we have tested both automatic differentiation and adjoint method to acquire the gradient of $\mathrm{d}L/\mathrm{d}\theta_{s}$ and achieved comparable results. For simplicity and clarity, we only show the results from the automatic differentiation approach, where we have developed a numerical solver based on JAX-CFD [68, 69] to facilitate the extraction of gradient information.
@@
Figure 4 shows the evolution of the posterior distribution for $\{\theta_{x},\theta_{y}\}$ across five stages. The high probability regions gradually converge to the true source location, suggesting an effective correction of parametric model error $\theta_{s}$. The comparison between Figs 3 and 4 confirms that the proposed hybrid approach achieves similar results to standard BED for the joint distribution of $\{\theta_{x},\theta_{y},\theta_{s}\}$ but with lower computational cost. Specifically, in the standard full approach, the posterior is approximated over a discretized grid of size $51\times 51\times 31$, requiring 80,631 forward model evaluations per stage. In contrast, the proposed hybrid approach only requires $51\times 51=2601$ evaluations per stage, leaving the task of parameter estimation for $\theta_{s}$ to the gradient-based optimization as described in Section 2.2. It is worth noting that the proposed hybrid approach has more computational advantage than the full Bayesian approach when structural model error exists, for which the model error needs to be parametrized by higher-dimensional unknown parameters and we provide a comprehensive study in Section 3.2.
@@
Figure 4: Posterior results of BED parameter $\{\theta_{x},\theta_{y}\}$ via the hybrid approach on parametric error. The $\{\theta_{x},\theta_{y}\}$ space is $0,1]^{2}$ as before, with a zoomed-in view $0.2,0.6]\times[0.1,0.5]$ to highlight detailed behaviors.
@@
We further investigate the results of corrected parametric model error $\theta_{s}$. As shown in Fig. 5(a), the proposed hybrid approach identifies the true value of $\theta^{\dagger}_{s}=2$. However, the estimate of $\theta_{s}$ tends to converge towards an incorrect value between iterations 450 and 600 (stage 4) during the optimization process. This deviation highlights the impact of experimental design quality on model error correction. More specifically, different designs yield varying data informativeness for $\theta_{s}$: some lead to robust estimation toward the true value, while other introduces bias. It is worth noting that the data quality cannot be quantified with the amount of biases in either the corrected model $\mathcal{G}+\mathrm{NN}$ or the simulated system states, as the truth of both is unknown a priori. To address this issue, we employ the approximated KLD defined in Eq. (18) to practically evaluate the data informativeness. Specifically, the original design location at stage 4 is $\{0.44,0.32\}$, which we refer to as the poor design due to its limited ability to reduce the bias in $\theta_{s}$. On the other hand, we randomly check several nearby points and find that better calibration results of $\theta_{s}$ can be achieved by a nearby alternative design point $\{0.41,0.26\}$, which is considered to yield more informative data and is referred to as the good design. In Fig. 5(b), we illustrate how $\theta_{s}$ evolves under the two different designs by plotting the ensemble mean and variance at each EKI iteration. The poor-quality design at stage 4 not only renders a bias in the corrected value for the parametric model error $\theta_{s}$ but also leads to a slower rate of variance reduction during EKI iterations, which implies a smaller approximated KLD between the prior and posterior ensemble of $\theta_{s}$. The comparison in Fig. 5(b) confirms that the performance of the ensemble-based approximated KLD as an indicator to evaluate data informativeness for the calibration of parametric model error.
@@ 3.2 Hybrid Approach on Structural Errors
which is different from the exponentially decaying source term as defined in Eq. (21) for the true system. Without addressing this model structural error, the inference of source location through standard BED yields biased results. In this work, a neural network $\text{NN}(z_{x}-\theta_{x},z_{y}-\theta_{y};\bm{\theta}_{\text{NN}})$ is employed to characterize the model structural error. More specifically, we utilize a fully connected neural network consisting of 37 parameters, a configuration that introduces a higher-dimensional parameter space than the example of parametric model error in Section 3.1. This network takes a 2-dimensional input vector, representing the relative distance between a given grid point and the source location, to output a scalar correction at each spatial point. This output quantifies the difference between the true and modeled source term values. The architecture features two hidden layers, each with four neurons and a tanh activation function. Since the governing equation’s RHS source term follows an inherently pointwise exponential or reciprocal decay function, a fully connected feedforward neural network mapping individual points independently is a natural choice. Since our coordinate‐based network is trained only on a limited set of pointwise data from the BED framework and optimized without any gradient‐ or residual‐based loss terms, it cannot reliably capture the non-local spatial information for characterizing convection and diffusion. On the other hand, the neural network output is time-invariant, while the convective velocity increases over time, it is incapable of matching the full RHS alone and thus cannot replace the role of existing physics-based terms in the modeled system. In addition, a structural constraint is imposed such that the neural network approximates a spatially decaying function that centers at its input reference point, i.e., the estimated source location $\{\theta_{x},\theta_{y}\}$, which is the same center as the physical-based source term in the RHS of the governing equation. In our numerical example, this constraint is incorporated via a pre-training phase, i.e., the network is initialized on data generated from the true target decaying source form. It is worth noting that such pre-training data would not guarantee the desired constraint in the subsequently trained model and may not be available in real-world applications. An alternative approach to enforce desired constraints on the trained neural network model is to add a penalty term into the loss function, e.g., penalizing the positive gradient with respect to the outward direction from the reference center point in this numerical example, to ensure the trained neural network is a spatially decaying function.
@@
We employ a uniform prior distribution over the range $0,1]^{2}$ for $\{\theta_{x},\theta_{y}\}$, discretized into a $51\times 51$ grid, with the true value of the source location set at $\{0.25,0.25\}$. The values of $\{\theta_{s}^{\dagger},\theta_{h}^{\dagger}\}$ are set as $\{2,0.05\}$ in both the true system and the modeled one. The advection velocity is known and set as $v_{x}=v_{y}=50t$. Design constraint remains the $0.2,0.2]^{2}$.
@@
We first present the results of standard sequential BED for physics-based parameters updated without correcting the structural error. It is expected that model discrepancy introduces bias in both the model forecast and likelihood values, resulting in an incorrect posterior distribution, i.e., the high probability area does not cover the true source location in Fig. 6(a). Fig. 6(b) shows that the posterior distribution can be improved by applying the proposed hybrid framework to correct the model error, and the high probability areas gradually converge to the true source location from stage 1 to stage 3. The deviation of high probability areas from the true source location in stages 4 and 5 is mainly caused by the uninformative design and data used for model error correction at stage 3. These design and data, optimized for physical parameters at stage 3, introduce biases into the calibrated model, which propagate into the Bayesian updating of physical parameters in subsequent stages. This observation does not contradict the Bayesian framework, as the posterior may still shift away from the prior if the likelihood is sharply concentrated in regions with non-negligible prior mass.
@@
Figure 7 shows that a more informative design and the corresponding measurement data (adopted at stage 3) for the model error correction significantly improve the posterior distributions of inferred source location at stages 4 and 5. We present the results using the ensemble-based approximated KLD to assess the data informativeness of a given design in Fig. 8. Two designs are selected for comparison: (i) the design identified by BED for posterior estimation at stage 3, represented by the orange dot in Fig. 7, and (ii) a modified design represented by the blue triangle in the same panel. With the untrained model at stage 3 as a baseline, we first generate an ensemble of its parameters and apply EKI to update the ensemble for 10 iterations using the two designs and corresponding measurement data. The ensemble-based approximated KLD between the updated and initial ensembles at each EKI iteration in Fig. 8(b) shows a faster increase with the good design, which confirms that the ensemble-based indicator proposed in Section 2.3 is effective of assessing data informativeness. The model error updated by the good design also shows less deviation from the true discrepancy in Fig. 8(a). Considering that the measurement data at a few design points are not sufficient to correct the model error term across the whole domain, some remaining differences in Fig. 8(a) between the true source term and the modeled one are expected. Even with those differences in the modeled source term, especially at a smaller distance from the source location, the posterior distributions in Fig. 7 still demonstrate a good estimation of the inferred source location.
@@
MSE $=\frac{1}{N_{z}}\sum_{z_{x},z_{y}}(\mathbf{u}(z_{x},z_{y})-\mathbf{u}^{\dagger}(z_{x},z_{y}))^{2},$ (23)